OpenStudy (anonymous):

Let R=qs-4 .when s=8 andR=16.when s=10,R is equal to

OpenStudy (anonymous):

please reply this question

OpenStudy (anonymous):

get q using the s and R

OpenStudy (anonymous):

s=8 R=16

OpenStudy (anonymous):

get q

OpenStudy (anonymous):

the answer is 21

OpenStudy (anonymous):

16=8q-4 20=8q q=20/8

OpenStudy (anonymous):

no .the correct answer is 21

OpenStudy (anonymous):

we are not done so R=20/8(s)-4 R=20/8(10)-4=25-4=21

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

272/3multiply 81-1/2 =3x what is the vaue of x